OpenStudy (anonymous):

solve the formula v= pi radius^2h for h

OpenStudy (amistre64):

your 'h' looks deceptive. But since its for the area of a circle; I assume its on it own perhaps?

OpenStudy (anonymous):

h=v/(pi*r^2)

OpenStudy (amistre64):

or it could simply be a random equation with an exponent of 2h

OpenStudy (anonymous):

i guess v is the volume and h is the height...a cylinder

OpenStudy (amistre64):

... that does match pretty well :)

OpenStudy (amistre64):

\(v=\pi r^2 h\) ; divide both sides by \(\pi r^2\) \[\frac{v}{\pi r^2}=h\]